show that the series ∑∞n=1 sin n theta/n does not converge uniformly on the interval ]0,2π[.


1
Expert's answer
2021-06-15T16:07:05-0400

"\sum_{n=1}^{\infty} \frac{sin n\theta}{n}
, \theta\in]0,2\pi[.\\
Then,\\
f_n=\frac{sin n\theta}{n}
\to0as n\to \infty\\
f'_n=cosn\theta\\
f'_n(\pi)=(-1)^n\text{ does not converge as } n\to \infty
\\
\text{Thus, the given series is not uniformly covergent.}\\"